Perpetual Inventory Using FIFO Beginning inventory, purchases, and sales for Item ER27 are as follows: May 1 Inventory 58 units @ $16 9 Sale 40 units 13 Purchase 60 units @ $17 28 Sale 22 units Assuming a perpetual inventory system and using the first-in, first-out (FIFO) method, determine (a) the cost of merchandise sold on May 28 and (b) the inventory on May 31. a. Cost of merchandise sold on May 28 $ b. Inventory on May 31 $
Answer #1
Quantity Price Cost
Beginning inventory 58 16 928
Purchase May 13 60 17 1020
Total 118 1948

Cost of merchandise sold on may 28 = (18*16+4*17) = $356

Inventory on may $31 = (1948-928-68) = $952